OpenStudy (anonymous):

Write a proportion for the following problem: The salesman’s commission on a sale of $520 is $50. At the same rate, what is the commission on a sale of $942?

OpenStudy (anonymous):

\[\frac{50}{520}\cdot 942\]

OpenStudy (anonymous):

\[\frac{50}{520}\cdot\frac{942}{1}\] Complete this problem and then write out the decimal representation of the fraction.